Circular No. 6939

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980613 E. Costa, Istituto di Astrofisica Spaziale, CNR, Rome; L. A. Antonelli,

Osservatorio Astronomico di Roma and BeppoSAX, Rome; C. De Libero, G. Spoliti, M. Celidonio, A. Tesseri, and M. R. Daniele, BeppoSAX, Rome; J. in ’t Zand, Space Research Organization Netherlands, Utrecht; and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report: “The error circle of GRB 980613 (IAUC 6938) was observed with the BeppoSAX Narrow Field Instruments from June 13.5618 (9 hr after the burst) to 14.6771 UT. Two previously unknown sources have been detected by the Medium Energy Concentrator Spectrometer (MECS) within the circle during this follow-up observation. One source, 1SAX J1018.1+7133, is located at α = 10^{h}18^{m}04^{s}, δ = +71^{o}33′30′′ (equinox 2000.0; error radius 50′′), and the other source, 1SAX J1017.9+7127, is located at α = 10^{h}17^{m}53^{s}, δ = +71^{o}27′24′′. 1SAX J1018.1+7133 has an average count rate of 0.0036 ± 0.0004 counts/s in the two MECS units, corresponding to a flux of (1.9 ± 0.2) × 10^{−13} erg cm^{−2} s^{−1}. This source does not show substantial variability. 1SAX J1017.9+7127 has an average count rate of 0.0016 ± 0.0004 counts/s in the two MECS units, corresponding to a flux of (1.1 ± 0.3) × 10^{−13} erg cm^{−2} s^{−1}. In the first 7000 s, the average count rate is ∼ 3 times this value; due to this fading behavior, the source is likely to be the x-ray afterglow of GRB 980613.”

MICROLENSING EVENT D. Bennett, University of Notre Dame, reports for the MACHO Col-

laboration (cf. IAUC 6845) plus affiliate S. Rhie: “Further observations of microlensing event MACHO-98-SMC-1 (cf. IAUC 6935) with the Mount Stromlo 1.3-m MACHO survey telescope and the Cerro Tololo 0.9-m telescope confirm the binary-lens interpretation and yield a prediction for the time of the second caustic crossing of June 19.2 ± 1.5 UT. Photometric measurements are urgently requested because the crossing timescale will indicate whether this binary lens resides in the Small Magellanic Cloud (with the star then expected to reach R ∼ 18 at the peak of a ∼ 12–24-hr crossing) or in the Milky Way’s dark halo (with expected R < 17 at the peak of a ∼ 1-hr crossing). Potential observers are urged to contact the MACHO team (see IAUC 6935) to coordinate observations. For additional information and

updates on this event, see http://darkstar.astro.washington.edu/98-
SMC-1 lev2.html and http://bustard.phys.nd.edu/MPS/98-SMC-1/.”

1998 June 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green